undeadknight26 (undeadknight26):

A system of equations is shown below: 6x – 5y = 5 3x + 5y = 4 The x-coordinate of the solution to this system of equations is _____.

OpenStudy (anonymous):

well you can multiply the top equation by -1 then you'll have: 5y-6x=-5 5y+3x=4 equate them: 6x-5=4-3x solve for 'x'

undeadknight26 (undeadknight26):

@PixieDust1 @AndyEmo @Mashy @cambrige @blondestarsbarbie @Nuke_ur_friend @ns_69 @sarahusher

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

after solving the way @sarahusher mentioned!!

undeadknight26 (undeadknight26):

Thank you both...sara would u mind medaling camb?

OpenStudy (anonymous):

Wait, when I solved it for the first time I got x=1/3 @sarahusher